# Aarti Ghatkesar

### v2

@@ Abstract
Achieving deep alignment between vision and language remains a central challenge for Multimodal Large Language Models (MLLMs). These models often fail to fully leverage visual input, defaulting to strong language priors. Our approach first provides insights into how MLLMs internally build visual understanding of image regions and then introduces techniques to amplify this capability. We explore techniques designed both to deepen the model’s understanding of visual content and to ensure that these visual insights actively guide language generation. We demonstrate the superior multimodal understanding of our resultant model through a detailed upstream analysis on predicting visually-dependent tokens as well as $>$10 percentage point boost on a visually challenging task and a consistent boost across multiple tasks.
@@ 1 Introduction
To address these fundamental challenges, this paper proposes a novel methodology centered on enhancing the model’s internal visual representation. First, we strengthen the visual signal processing, enabling the model to extract richer and more fine-grained visual representations. Second, we introduce a training mechanism that explicitly encourages the model to allocate greater attention to these enhanced visual inputs during the response generation phase, reducing over-reliance on language priors. Finally, complementing these training modifications, we construct a targeted synthetic dataset specifically designed to leverage this advanced training fabric, providing controlled examples that force the model to learn and exploit fine-grained visual cues effectively. Our key contributions are:
Enhanced Visual Representation using VisualLoss . We introduce a new loss function to ensure the LLM backbone builds a rich representation of the input image, independent of whether they are mentioned in associated training text.
@@ 3.2 Visual Representation using VisualLoss
Figure 2: Proposed modifications to MLLM training that enhances visual understanding through VisualLoss  (Section 3.2) and encourages LLM backbone to pay greater attention to them using BlankTokens  (Section 3.3).
@@ 4 Results
This section analyzes the effectiveness of our proposed approach. We layer our techniques on top of a LlaVA [9] model with Llama 3.1 8B backbone (Baseline) and detail the training setup in Appendix 6.4. We demonstrate that our approach improves the model’s ability to predict visually relevant tokens (Section 4.1) and significantly enhances its performance on challenging visual tasks (Section 4.2).
@@ 4.1 Upstream Analysis
To evaluate the impact of our techniques on the model’s ability to integrate visual information during language generation, we conduct an upstream analysis on the SpatialMM dataset [15], that by construction has text with a high dependence on specific visual content, demanding strong visual grounding, making it suitable for this analysis.
@@ 4.2 Evaluation on SpatialEval Benchmark
We evaluate our proposed techniques on the SpatialEval [19] benchmark111we had to enhance benchmark’s prompting and parsing capabilities which we open source as well for reproducibility, designed to probe spatial understanding capabilities. Our results, summarized in Table 2, demonstrate a clear trend of performance improvement across all benchmark subsets with our proposal.
@@
Technique Grid MazeNav Map Real Avg. Baseline 34.5 15.1 44.1 50.4 1 + VisualLoss 31.3 17.0 50.7 56.3 1.075 + BlankTokens 30.1 22.9 47.1 52.6 1.125 + Synthetic 41.1 26.3 55.2 53.3 1.31 + VisualLossAdv 43.5 26.7 64.1 53.3 1.385 Llama 3.2 11B [5] 50.1 28.1 48.2 53.3 1.365
4.3 Future Extensions
We explore two extensions to further improve LLM’s internal visual representation. First, to mitigate the potential representational conflict arising from tasking the same LLM weights with processing both visual and textual concepts, we experimented with providing each modality with an independent set of weights [20, 8]. Cross-modal interaction was maintained through the standard self-attention mechanism. Our initial results in Table 1 (+ind. weights row) indicate that this separation significantly enhances visual representation - providing the biggest reduction in upstream loss thus far ($3.48$Rightarrow $3.20$).
Next, we demonstrate that VisualLoss  can leverage advancements in image encoders to help the LLM backbone build richer visual representations. We replace I-JEPA [2] with the AM-Radio [12] encoder as the auxiliary image encoder for our visual loss (shown as VisualLossAdv  in Table 2) yields an additional gain of approximately 9.1 percentage points on the SpatialMap subset. This enhancement enables our model to match the performance of Llama 3.2 11B on this task, despite our model being smaller, employing a simpler architecture, and using a lower-resolution visual input. Further investigation of these extensions constitute key directions for our subsequent research.

### v3

# Perceiving Beyond Language Priors: Enhancing Visual Comprehension and Attention in Multimodal Models

@@ Abstract
Achieving deep alignment between vision and language remains a central challenge for Multimodal Large Language Models (MLLMs). These models often fail to fully leverage visual input, defaulting to strong language priors. Our approach first provides insights into how MLLMs internally build visual understanding of image regions and then introduces techniques to amplify this capability. We explore techniques designed both to deepen the model’s understanding of visual content ( VisualLoss ) and to ensure that these visual insights actively guide language generation ( BlankTokens ). We demonstrate the superior multimodal understanding of our resultant model, PerceptLLM , through a detailed upstream analysis on predicting visually-dependent tokens as well as $>$10 percentage point boost on a visually challenging task and a consistent boost across multiple tasks. [^1] Accepted to the Second Workshop on Visual Concepts at CVPR 2025.
@@ 1 Introduction
To address these fundamental challenges, this paper proposes PerceptLLM , multimodal models trained with a novel methodology centered on enhancing the model’s internal visual representation. First, we strengthen the visual signal processing, enabling the model to extract richer and more fine-grained visual representations. Second, we introduce a training mechanism that explicitly encourages the model to allocate greater attention to these enhanced visual inputs during the response generation phase, reducing over-reliance on language priors. Finally, complementing these training modifications, we construct a targeted synthetic dataset specifically designed to leverage this advanced training fabric, providing controlled examples that force the model to learn and exploit fine-grained visual cues effectively. Our key contributions are:
Enhanced Visual Representation via Auxiliary Loss and Disentangled Architecture. We introduce a novel approach that combines an auxiliary visual loss ( VisualLoss ) with a key architectural modification. The loss function ensures the LLM backbone builds a rich representation of the entire input image, independent of associated text, while our architectural change decouples the processing pathways for image and text tokens using independent weights. This disentanglement allows the model to specialize in rich visual representation learning without conflicting with its autoregressive text generation objective.
@@ 3.2 Visual Representation using VisualLoss
Figure 2: Proposed modifications to MLLM training that enhances visual understanding through VisualLoss  (Section 3.2) and encourages LLM backbone to pay greater attention to them using BlankTokens  (Section 3.3). We also highlight the tension between VisualLoss  and next-token prediction that is addressed through independent pathways for vision and text (Section 3.4).
@@ 3.3 Language Priors using BlankTokens
3.4 Disentangling Visual Representation and Autoregressive Prediction
While the auxiliary VisualLoss  described previously proved effective in enhancing model’s visual representation (Section 4), we also observe that it introduces an inherent tension in the LLM backbone’s training objectives. The model is simultaneously tasked with two fundamentally different goals depending on the input modality. For text tokens, the objective is standard causal language modeling: each token’s representation is optimized to predict the next token in the sequence. This is an explicitly autoregressive task that relies on a natural left-to-right flow. In contrast, for image tokens, our visual loss encourages the model to learn a rich representation of the current patch, as images lack the same inherent sequential or autoregressive structure. Forcing a single set of Transformer weights to handle both these conflicting, modality-dependent objectives leads to the weights being pulled in different directions.
To resolve this conflict, we introduce a key architectural modification: we decouple the processing of image and text tokens within the LLM backbone by assigning them independent sets of weights [9, 21]. By creating this programmatic distinction, we effectively allow for two specialized pathways to emerge within the Transformer layers. The pathway for image tokens can optimize its parameters, specifically the self-attention parameters, for the non-autoregressive task of building rich spatial representations. Concurrently, the pathway for text tokens can retain its well-established autoregressive capabilities for fluent and coherent language generation. This disentanglement allows for distinct attention behaviors tailored to each modality, enabling a more effective and specialized learning process for both visual understanding and textual prediction.
As demonstrated empirically in our results (Section 4.2), this architectural modification yields a significant improvement in model accuracy, providing strong empirical validation for our hypothesis. Importantly, this approach maintains deployment efficiency. Despite the increase in the total number of parameters to support both modalities, the model incurs no additional computational or memory access overhead during inference, since only the pathway corresponding to the token’s modality (image or text) is utilized at any given step.
@@ 4 Results
This section analyzes the effectiveness of PerceptLLM . We layer our techniques on top of a LlaVA [10] model with Llama 3.1 8B backbone (Baseline) and detail the training setup in Appendix 6.4. We demonstrate that our approach improves the model’s ability to predict visually relevant tokens (Section 4.1) and significantly enhances its performance on challenging visual tasks (Section 4.2).
@@ 4.1 Upstream Analysis
To evaluate the impact of our techniques on a model’s ability to integrate visual information during language generation, we conduct an upstream analysis on the SpatialMM dataset [16], that by construction has text with a high dependence on specific visual content, demanding strong visual grounding, making it suitable for this analysis.
@@ 4.2 Evaluation on SpatialEval Benchmark
We evaluate our proposed techniques on the SpatialEval [20] benchmark222we had to enhance benchmark’s prompting and parsing capabilities which we open source as well for reproducibility, designed to probe spatial understanding capabilities. Our results, summarized in Table 2, demonstrate a clear trend of performance improvement across all benchmark subsets with our proposal.
@@
Building on these improvements, we next investigate if our framework can leverage advancements in image encoders to further enhance the LLM’s internal visual representation. We find that replacing I-JEPA [2] with the AM-Radio [13] encoder as the auxiliary image encoder for our visual loss (shown as VisualLossAdv  in Table 2) yields an additional gain of approximately 9.1 percentage points on the SpatialMap subset. This enhancement enables our model to match the performance of Llama 3.2 11B on this task, despite our model being smaller, employing a simpler architecture, and using a lower-resolution visual input.
Our final model, PerceptLLM , integrates all proposed enhancements - auxiliary visual loss ( VisualLoss ), input masking ( BlankTokens ), the disentangled architecture with independent pathways (Section 3.4), synthetic data (Section 3.5) and an advanced AIM-v2 image encoder [5]. This configuration yields our strongest results, significantly outperforming the larger Llama 3.2 11B model while being more efficient in terms of model size, architecture, and input resolution.
Future Work: An interesting aspect of PerceptLLM  design is it’s larger capacity and possibility of independently impacting pathways for vision and text. We expect that this departure in model size and capability vs the baseline model could lead to baseline data mixture and hyper-parameter configuration being sub-optimal for PerceptLLM . For instance, the model’s enhanced capacity may make it more susceptible to overfitting on certain peculiarities within the synthetic dataset. This suggests that further tuning of synthetic data generation as well as ablation studies on the data mixture for this model represent a promising avenue for future optimization.
@@
Technique Grid MazeNav Map Real Avg. Baseline 34.5 15.1 44.1 50.4 1 + VisualLoss 31.3 17.0 50.7 56.3 1.075 + BlankTokens 30.1 22.9 47.1 52.6 1.125 + Synthetic 41.1 26.3 55.2 53.3 1.31 + VisualLossAdv 43.5 26.7 64.1 53.3 1.385 PerceptLLM 50.8 35.0 55.13 52.6 1.52 Llama 3.2 11B [6] 50.1 28.1 48.2 53.3 1.365
@@ 5 Conclusion
Acknowledgement
Thanks to Shane Segal for leading the implementation of our disentangled architecture, adapting concepts from Mixture-of-Transformers [9] to create specialized processing pathways for the vision and text modalities. Thanks to Uddeshya Upadhyay for building the pipeline for synthetic data generation and feedback on the early drafts.